# A Forward Reachability Perspective on Control Barrier Functions and Discount Factors in Reachability Analysis

### v3

@@ 4.1 FRT Value Function with Discount Factor
The value function (18) captures a differential game between the control and the disturbance, wherein the optimal control signal of this game is verifying the existence of a trajectory that reaches $x$ without passing through $\mathrm{Int}(S)$ in the past under the worst-case disturbance. If such a trajectory does not exist, $V_{\gamma}(x)$ is positive and $x$ is inside $\mathrm{FRT}(\mathrm{Int}(S))$.
We now discuss the effect of introducing $\gamma$ to the cost function. (18) defines a differential game with a discounted supremum-over-time cost function. Larger $\gamma$ will recognize the value of $h_{S}$ at the current time more than the value in the past. More importantly, $\gamma$ would also affect the resulting optimal control policy; whereas when there is no discount, the optimal control always has to try its best to maintain the value of $h_{S}$, the discount alleviates this conservativeness and allows the optimal control to decay the value of $h_{S}$.
@@
As such, the discount factor introduces the “game-of-degree” aspect to the reachability problem. In this game-of-degree, the parameter $\gamma$ serves as a knob that adjusts how conservative the resulting optimal policy will be. However, the fundamental nature of the reachability problem, what is called the “game-of-kind” [12, 6]—whether or not a state is inside the FRT—remains consistent, since (20) holds for any value of $\gamma$ and the property can be determined by the positivity of $V_{\gamma}(x)$.
@@
The condition $L_{f}<\gamma$ implies that the discount factor has to be large enough to suppress the effect of the vector field in prohibiting continuity. Under this condition, since the value function is Lipschitz continuous, it is differentiable almost everywhere by Rademacher’s Theorem. As we will later see, other infinite-horizon value functions in backward reachability formulations [17, 13] do not have Lipschitz continuity and can even be discontinuous, which prohibits the usage of a differential inequality-based condition like the barrier constraint for deriving safe control policies from the value function.
The contraction property of the Bellman backup will be discussed next after introducing the HJ-PDE characterization of $V_{\gamma}$, which provides a computational machinery for the computation of the value function $V_{\gamma}$.
@@ 4.2 Hamilton-Jacobi Characterization
Suppose $h_{S}$ is a bounded and Lipschitz continuous function, and $\gamma>0$. $V_{\gamma}$ in (18) is a unique viscosity solution [8] in $\mathbb{R}^{n}$ of the following HJ-PDE, called forward reachable tube Hamilton-Jacobi variational inequality (FRT-HJ-VI):
[…]
$\max_{u}\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\Big{\}}.\vspace{-0.5em}$
@@
For non-positive values of $\gamma$, $V_{\gamma}$ might be unbounded and the FRT-HJ-VI might have multiple solutions. In contrast, a strictly positive value of $\gamma$ guarantees the boundedness and the uniqueness of the solution of the FRT-HJ-VI. (An example in Appendix illustrates these outcomes.) In fact, the uniqueness property follows from the contraction property of the Bellman backup associated with the dynamic programming principle of $V_{\gamma}$.
@@
[…]
$\!\!\!e^{-\gamma T}V(\mathrm{x}^{-}(-T))\Big{\}}.\vspace{-1em}$ (22)
@@
Theorem 3 (Contraction mapping).
@@
$\lVert B_{T}[V^{1}]-B_{T}[V^{2}]\rVert_{\infty}\leq e^{-\gamma T}\lVert V^{1}-V^{2}\rVert_{\infty},\vspace{-0.75em}$ (23)
and the FRT value function $V_{\gamma}$ in (18) is the unique fixed-point solution to $V_{\gamma}\!=\!B_{T}[V_{\gamma}]$ for each $T\!>\!0$. Also, for any $V\!\in\!\textnormal{BUC}(\mathbb{R}^{n})$, $\lim_{T\rightarrow\infty}B_{T}[V]=V_{\gamma}$.
@@
Theorem 3 provides various ways to compute the FRT value function $V_{\gamma}$ using the operation $B_{T}[\cdot]$, which do not require any assumptions for the initial guess of the value function, besides the boundedness and uniform continuity in $\mathbb{R}^{n}$. For instance, it allows the computation of $V_{\gamma}$ with a finite-horizon HJ-PDE, for which numerical solution methods like the level set method [29] are well established. The finite-horizon value function is guaranteed to converge to $V_{\gamma}$ as $T\rightarrow\infty$. Also, the theorem enables other numerical schemes based on time discretization, such as value iteration. For further details of how Theorem 3 can be utilized for finite-horizon-based computation or value iteration, see Appendix 7.5.
4.3 FRT Value Functions for Control Invariant Sets
@@
First, when $S$ has a differentiable boundary and is robustly control invariant where Theorem 1 holds, the following holds:
Proposition 3 (FRT value function characterization of robust control invariant sets).
Under Assumptions 2, 3, and 4, $S$ is robustly control invariant if and only if
$\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)\;$ =$=\{x~{}|~{}V_{\gamma}(x)>0\},$
$\mathrm{Int}(S)^{c}$ =$=\{x~{}|~{}V_{\gamma}(x)=0\},$
where $V_{\gamma}$ is defined as (18).
Proof.
This results from Theorem 1 and Lemma 7. ∎
When Proposition 3 holds, the control invariant set $\mathrm{Int}(S)$ is characterized as a strict zero-superlevel of $V_{\gamma}$. This enables the synthesis of a control policy using $V_{\gamma}$ to maintain forward invariance of trajectories within $\mathrm{Int}(S)$. To see this, we derive an optimal policy of $V_{\gamma}$ from the FRT-HJ-VI (21).
@@
where $V_{\gamma}$ is defined as (18). Then, $K_{\gamma}(x)$ is non-empty for every $x\in\mathrm{Int}(S)$ where $\frac{\partial V_{\gamma}}{\partial x}$ exists. In addition, if $V_{\gamma}$ is differentiable, any element of $K_{\gamma}(x)$ is an optimal control input with respect to $V_{\gamma}$ in (18), and under Assumptions 2, 3, and 4, if $S$ is robustly control invariant, the trajectory under $K_{\gamma}(x)$ remains forward invariant in $S$ under the worst-case disturbance.
@@
Note that the non-emptiness of $K_{\gamma}(x)$ is derived from $V_{\gamma}$ satisfying the FRT-HJ-VI (21). By noting that the second term of the minimum in (21) has to be non-negative for (21) to hold, we get that
$\max_{u\in U}\min_{d\in D}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\geq 0\vspace{-0.5em}$ (25)
@@
Since the value function is Lipschitz continuous and differentiable almost everywhere by Proposition 2, $V_{\gamma}$ satisfies (25) almost everywhere in $\mathrm{Int}(S)\subset\mathbb{R}^{n}$. Note that $V_{\gamma}$ is 0 everywhere outside $\mathrm{Int}(S)$. If $V_{\gamma}$ is differentiable in $\mathrm{Int}(S)$, (25) is satisfied everywhere in $\mathrm{Int}(S)$, which constitutes the robust CBF in Definition 4:
@@
If $V_{\gamma}$ is continuously differentiable in $\mathrm{Int}(S)$, $V_{\gamma}:S\rightarrow\mathbb{R}$ is a robust CBF.
Proof.
Proposition 4 implies that the barrier constraint holds at any state in $\mathrm{Int}(S)$. As we constrain the domain of $V_{\gamma}$ to $S$, the gradient of $V_{\gamma}$ at $x\in\partial S$, is defined as $\lim_{y\rightarrow x}\frac{\partial V_{\gamma}}{\partial x}$. Since $V_{\gamma}$ is continuously differentiable, this limit exists and $K_{\gamma}(x)$ in (24) is nonempty by Proposition 4. Thus, the statement holds by the definition of robust CBF in Def. 4. ∎
More importantly, any valid robust CBF $h$ itself is the FRT value function in $\mathrm{Int}(S)$:
@@
Under Assumptions 2 and 4, let $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a differentiable function that satisfies Assumption 3 and is a robust CBF for a closed set $S$, satisfying
$\max_{u\in U}\min_{d\in D}\frac{\partial h}{\partial x}\cdot f(x,u,d)+\gamma h(x)\geq 0,\vspace{-0.75em}$ (26)
@@
$V_{\gamma}(x)=\max\{0,h(x)\}\vspace{-0.75em}$ (27)
is the unique viscosity solution of the FRT-HJ-VI (21) with $h_{S}(x)=h(x)$.
@@
4.4 Example: Pendulum
We present an example of a pendulum system subjected to disturbance where we demonstrate the robustness of the safety control derived from Proposition 4. We use Proposition 4 to design the following robust safety filter:
Robust min-norm safety filter
$\pi_{S}(x,t)=\arg\min_{u\in U}\left\lVert u-u_{ref}(t)\right\rVert$ (28a)
s.t. $\min_{d\in D}\frac{\partial h}{\partial x}\cdot f(x,u,d)+\gamma h(x)\geq 0,\vspace{-1em}$ (28b)
where $h$ is the chosen CBF. The controller (28) filters a reference control signal $u_{ref}(t)$—when $u_{ref}(t)$ does not satisfy the barrier constraint, it selects a control input $u\in U$ that is closest to $u_{ref}(t)$ that satisfies (28b). Note that from Proposition 4, if we use an FRT value function $V_{\gamma}$ that is differentiable for $h$, the filter is always feasible for $x\in\mathrm{Int}(S)$, and will render the trajectory forward invariant in $\mathrm{Int}(S)$. In the case when the system is affine in control input and disturbance, this safety filter can be implemented as a quadratic program [4, 13].
The dynamics of the pendulum system is given as
$\begin{bmatrix}\dot{x_{1}}\\ \dot{x_{2}}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -\sin{x_{1}}\end{bmatrix}+\begin{bmatrix}0\\ 1\end{bmatrix}u+\begin{bmatrix}0\\ \cos{x_{1}}\end{bmatrix}d,\vspace{-0.5em}$ (29)
where $x_{1}=\theta$, $x_{2}=\dot{\theta}$ are the angle and the angular rate of the pendulum, respectively, $u$ is the applied torque, the control input, and $d$ is the horizontal acceleration applied to the base of the pendulum, the disturbance to the system. $x_{1}=0$ at the released configuration. The desired safety constraint is $X=\{x~{}|~{}0.5\pi\leq x_{1}\leq 2\pi,|x_{2}|\leq 1\}$, constraining both the range of the angle and the angular rate. We set the maximum torque $\bar{u}=\sin{\frac{\pi}{3}}$, so that the maximum torque cannot resist the torque produced by the gravity in the range of $\theta\in[\frac{\pi}{3},\frac{2\pi}{3}]$ and $\theta\in[\frac{4\pi}{3},\frac{5\pi}{3}]$. Then $U$ is set as $[-\bar{u},\bar{u}]$, and $D$ is set as $[-0.1,0.1]$.
Figure 3: (a) The desired safety constraint for the pendulum example is set to be $X=\{x|0.5\pi\leq\theta\leq 2\pi,|\dot{\theta}|\leq 1\}$. The robust control invariant set $S$ is designed by a Bezier fitting to the maximal control invariant set of $X$, computed from the backward reachability analysis. (b) The corresponding target function $h_{S}(X)$ and the computed FRT value function $V_{\gamma}(x)$ with $\gamma=5$.
The reference control is produced by a clipped feedback linearization controller $u_{ref}(t)=\pi_{ref}(\mathrm{x}(t))$, given as
$\pi_{ref}(x)\!=\!\min\!\left\{\max\{\sin{x_{1}}\!-\!k_{1}(x_{1}\!-\!x_{1,\text{t}})\!-\!k_{2}x_{2},-\bar{u}\},\bar{u}\right\},\vspace{-0.5em}$
where the desired target angle $x_{1,\text{t}}$ is set as $-0.2$ for $t\in[0,8]$ and $\pi-0.6$ for $t>8$.
For the design of the target function $h_{S}$, we first verify the maximal control invariant set contained in $X$, using the backward reachability analysis, presented in [17, 38]. Then, a control invariant set with a differentiable boundary, $S$, and its distance function $h_{S}$, is designed by applying Bezier curve fitting [30] to the maximal control invariant set. The resulting $S$ and $h_{S}$ are visualized in Figure 3. We then compute the FRT value function $V_{\gamma}$. The result when $\gamma=5$ is presented in Figure 3, which is differentiable in $\mathrm{Int}(S)$.

### v4

@@ 3.2 Forward Reachable Tubes of Control Invariant Sets
Condition 1 is crucial for the theorem’s proof, which establishes the equivalence between $f(x,\bar{\pi}(x))\in T_{S}(x)$ and $f(x,\bar{\pi}(x))\in T_{\mathrm{Int}(S)^{c}}(x)$ (Lemma 5), where $S=\mathrm{FRT}(\mathrm{Int}(C))$. At the non-differentiable boundary of a set $S$, this is not necessarily true. In this case, a forward trajectory can be inevitably “leaked” from the interior of the FRT through the point at the boundary where it is non-differentiable. An example of this incident is introduced in Section 3.3 (Figure 2a).
In addition, the following corollary shows that the interior of any control invariant set with a differentiable boundary is a fixed point of the $\mathrm{FRT}(\cdot)$ set operator.
Corollary 1.
A set $S$ satisfying Condition 1 is robustly control invariant under (8) if and only if $\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)$.
Proof.
See Appendix 8.2. ∎
As noted in Remark 1, since control invariant sets associated with any CBFs have to satisfy Condition 1, Corollary 1 implies that this fixed point property is a key feature of a valid CBF design.
Finally, in the case when the FRT or the initial set $C$ does not satisfy Condition 1, the FRT stills provide a valuable information in characterizing an invariant set that contains $C$:
Theorem 2.
For any given set $C$, the minimal robust control invariant set $S$ which includes $C$ and satisfies Condition 1 is always a superset of $\mathrm{FRT}(\mathrm{Int}(C))$.
Proof.
See Appendix 8.3. ∎
Note that again, Condition 1 is necessary for the invariant set $S$ to be the superset of $\mathrm{FRT}(\mathrm{Int}(C))$. A counterexample when Condition 1 does not hold is explained in Section 3.3 (Figure 2b).
This theorem implies that the FRT can be used to lower-bound when finding a control invariant set that contains $C$. We extend this property further in Sections 4.3 and 5 to devise a CBF design method whose control invariant set overappoximates the FRT.
3.3 Example: Double Integrator
We introduce examples of four different initial sets (Figure 2 first row) and their FRTs (Figure 2 second row) to illustrate the importance of the differentiability of its boundary for control invariance. These examples are based on a simple double integrator system defined by $\dot{p}=v,\dot{v}=u$, with state $x=[p\;v]^{T}$, and control input $u$, with control bound $u\in[-1,1]$. Note that curves $p=c\pm\frac{1}{2}v^{2}$ characterize trajectories with the saturated input $u=\pm 1$, $\ddot{p}=\pm 1$, that decelerate or accelerate until $v=0$, . The four specific $C$ sets are:
  1. (a)
    $C_{a}$ is defined as the circular region with radius $r$ centered at the origin. This set $C_{a}$ satisfies Condition 1 but is not control invariant.
  2. (b)
    $C_{b}$ is formed by five curves, $p_{1}+\frac{1}{2}v^{2}, p=p$$_{1}-\frac{1}{2}v^{2}, p=p$$_{2}-\frac{1}{2}v^{2}, p=p$$p_{2}+\frac{1}{2}v^{2}$, $p=-\frac{1}{2}v^{2}$, and the $p$-axis, as shown in Figure 2(b). This set does not satisfy Condition 1, but is control invariant.
  3. (c)
    $C_{c}$ is formed by two curves, $p=-p_{1}+\frac{1}{2}v^{2}$, and $p=p_{1}-\frac{1}{2}v^{2}$. This set also does not satisfy Condition 1 but is control invariant.
  4. (d)
    $C_{d}$ is formed by two curves, $p=-p_{1}+\frac{1}{2}v^{2}$ and $p=p_{1}-\frac{1}{2}v^{2}$ ($p_{1}>1$), and two arcs whose radius is $r=-1+2\sqrt{p_{1}}$ that are tangential to the curves whose centers are positioned at $(-p_{1}+r,0),(p_{1}-r,0)$, respectively. This set satisfies Condition 1 and is also control invariant.
Figure 3: Minimal control invariant set that contains $S_{a}$ in Figure 2a, which is a superset of $\mathrm{FRT}(\mathrm{Int}(C_{a}))$.
The first case demonstrates that the FRT can be a strict superset of $\mathrm{Int}(C_{a})$ when $C_{a}$ is not control invariant. The resulting $\mathrm{FRT}(\mathrm{Int}(C_{a}))$ is still not control invariant since the trajectory is bound to escape the set at points A and B. The actual minimal control invariant set that contains $C_{a}$ is visualized in Figure  3. This example reveals a challenge in constructing a control invariant set with exact FRT, when the initial set is not control invariant. Thus, in the next section, we focus on utilizing Theorem 2 to find an outer-approximated FRT that is control invariant.
In the second and third cases, the control invariance of $C_{a}$ and $C_{b}$ can be checked analytically. The second case shows that the FRT can be a strict superset of $\mathrm{Int}(C_{b})$ if Condition 1 is not met, even though $C_{b}$ is control invariant, showing the necessity of Condition 1 in Corollary 1. Note that point C is where $f(x,\bar{\pi}(x))\!\in\!T_{C}(x)$ holds but $f(x,\bar{\pi}(x))\!\in\!T_{\mathrm{Int}(C)^{c}}(x)$ does not hold. Moreover, since $C_{b}$ itself is the minimal control invariant set containing $C_{b}$, Theorem 2 does not hold due to the absence of Condition 1.
In the third case, since $f(x,\bar{\pi}(x))\!\in\!T_{\mathrm{Int}(C)^{c}}(x)$ holds at both points D and E where $\partial C_{c}$ is not smooth, $\mathrm{FRT}(\mathrm{Int}(C_{c}))=\mathrm{Int}(C_{c})$. Finally, the set $C_{d}$ in the last case satisfies Condition 1 and is also control invariant. Thus, according to Corollary 1, $\mathrm{FRT}(\mathrm{Int}(C_{d}))$ remains the same as $\mathrm{Int}(C_{d})$.
4 FRT value function and CBF
In this section, by taking the Hamilton-Jacobi approach to the forward reachability problem, we pose the computation of FRT as a differential game. The FRT is characterized by the value function proposed in Section 4.1. This value function is the unique solution to the HJ-PDE proposed in Section 4.2. In Section 4.3, we establish a connection between the proposed value function and the CBFs, by taking the supersolution of the HJ-PDE. This provides an interpretation of any valid CBF as a forward reachability value function, and also a constructive way to design CBF as a supersolution of the FRT value function.
4.1 FRT Value Function with Discount Factor
By noting that $h_{C}$ satisfying Condition 2 serves as a distance-like metric to the boundary of $C$ and its sign serves as an indicator of the inclusion in $C$, we can rewrite the definition of the FRT in (17) as follows:
$\mathrm{FRT}(\mathrm{Int}(C))$
$=\!\big\{x~|~\exists\xi_{d}^{-}\in\Xi_{d}^{-},\forall\mathrm{u}^{-}\!\in\!\mathcal{U}^{-},\!\sup_{t\in(-\infty,0]}h_{C}(\mathrm{x}^{-}(t))\!>\!0\big\}$
$=\!\big\{x~|~\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}h_{C}(\mathrm{x}^{-}(t))>0\big\}.\vskip-7.5pt$
Since rescaling $h_{C}(\mathrm{x}^{-}(t))$ with a positive constant at any time $t$ does not change its sign, the following holds:
$\mathrm{FRT}(\mathrm{Int}(C))\!=\Big\{x~|\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}e^{\gamma t}h_{C}(\mathrm{x}^{-}(t))>0\Big\},$
where $\gamma\!>\!0$ and at each time $t\!\in\!(-\infty,0]$, $h_{C}(\mathrm{x}^{-}(t))$ is rescaled by $e^{\gamma t}$. Thus, by defining the FRT value function of $C$, $V_{\gamma}:\!\mathbb{R}^{n}\!\rightarrow\!\mathbb{R}$, as
$V_{\gamma}(x)\coloneqq\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}J_{\gamma}(x,\mathrm{u}^{-},\Xi_{d}^{-})\vskip-10.00002pt$ (18)
$J_{\gamma}(x,\mathrm{u}^{-},\xi_{d}^{-})=\sup_{t\in(-\infty,0]}e^{\gamma t}h_{C}(\mathrm{x}^{-}(t)),\vskip-10.00002pt$ (19)
where $\mathrm{x}^{-}$ solves (16) and $x$ is the terminal state of $\mathrm{x}^{-}$.
@@ 4.1 FRT Value Function with Discount Factor
The value function $V_{\gamma}$ captures a differential game between the control and the disturbance, wherein the optimal control signal of this game is verifying the existence of a trajectory that reaches $x$ without passing through $\mathrm{Int}(C)$ in the past under the worst-case disturbance. If such a trajectory does not exist, $V_{\gamma}(x)$ is positive and $x$ is inside $\mathrm{FRT}(\mathrm{Int}(C))$, described by the following lemma.
Lemma 4.
Suppose $C\in\mathbb{R}^{n}$ is a closed set, and a bounded function $h_{C}$ satisfies Condition 2-(1). $V_{\gamma}(x)$ is positive if and only if $x$ belongs to the FRT of the interior of $C$:
$\mathrm{FRT}(\mathrm{Int}(C))$ $=\{x~|~V_{\gamma}(x)>0\},$
$\{\mathrm{FRT}(\mathrm{Int}(C))\}^{c}$ $=\{x~|~V_{\gamma}(x)=0\}.$ (20)
Proof.
See Appendix 8.4. ∎
As in Lemma 4, $\mathrm{FRT}(\mathrm{Int}(C))$ is characterized as the strict zero-superlevel set of $V_{\gamma}$ in (18). If its boundary is differentiable (Condition 1), then the FRT becomes a control invariant set according to Theorem 1. In such cases, whereas the BRT value function (14) flattens to zero within the control invariant set, the FRT value function remains positive inside the invariant set.
@@
The fundamental nature of reachability, what is called the “game-of-kind” [12, 6]—whether or not a state is inside the FRT—remains consistent, since (20) holds for any value of $\gamma$ and the property is determined by the positivity of $V_{\gamma}(x)$. However, the discount factor $\gamma$ introduces the “game-of-degree” aspect [24] to the reachability problem . In this game-of-degree, $\gamma$ serves as a knob that adjusts how conservative the resulting optimal policy will be. As we will see soon, this knob is shaped by the barrier constraint of the CBFs. This is the unique feature of introducing discount in FRT value functions whereas optimal control policy of a discounted BRT value function is not affected by $\gamma$ within the invariant set due to its flatness [2, 49].
@@
The condition $L_{f}<\gamma$ implies that $\gamma$ has to be large enough to suppress the effect of the vector field in prohibiting continuity. Under this condition, since the value function is Lipschitz continuous, it is differentiable almost everywhere by Rademacher’s Theorem.
The contraction property of the Bellman backup will be discussed next after introducing the HJ-PDE characterization of $V_{\gamma}$, which provides a computational machinery for the computation of the value function $V_{\gamma}$, and also the uniqueness of the PDE solution.
@@ 4.2 Hamilton-Jacobi Characterization
Suppose $h_{S}$ is a bounded and Lipschitz continuous function, and $\gamma>0$. $V_{\gamma}$ in (18) is a unique viscosity solution in $\mathbb{R}^{n}$ of the following HJ-PDE, called Hamilton-Jacobi forward reachable tube variational inequality (HJ-FRT-VI):
$0=\min\Big\{$ $\ V_{\gamma}(x)-h_{C}(x),$ (21)
$\max_{u}\min_{d}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\Big\}.\vskip-5.0pt$
@@
A strictly positive value of $\gamma$ guarantees the boundedness and the uniqueness of the solution of the HJ-FRT-VI. (An example in Appendix illustrates these outcomes.) In fact, the uniqueness property follows from the contraction property of the Bellman backup associated with the dynamic programming principle of $V_{\gamma}$.
@@
$B_{T}[V](x)\!\coloneqq\!\!\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\!\!\max\Big\{$ $\!\max_{t\in[-T,0]}e^{\gamma t}h_{S}(\mathrm{x}^{-}(t)),$
$\!\!\!e^{-\gamma T}V(\mathrm{x}^{-}(-T))\Big\}.\vskip-12.50002pt$ (22)
@@
Theorem 4 (Contraction mapping & Uniqueness).
@@
$\lVert B_{T}[V^{1}]-B_{T}[V^{2}]\rVert_{\infty}\leq e^{-\gamma T}\lVert V^{1}-V^{2}\rVert_{\infty},\vskip-7.5pt$ (23)
and the FRT value function $V_{\gamma}$ in (18) is the unique fixed-point solution to $V_{\gamma}\!=\!B_{T}[V_{\gamma}]$ for each T>0T\!>\!0. Also, for any $V\!\in\!\textnormal{BUC}(\mathbb{R}^{n})$, $\lim_{T\rightarrow\infty}B_{T}[V]=V_{\gamma}$.
@@
Theorem 4 also provides various ways to compute the FRT value function $V_{\gamma}$ using the operation $B_{T}[\cdot]$, which do not require any assumptions for the initial guess of the value function, besides the boundedness and uniform continuity in $\mathbb{R}^{n}$. For further details of how Theorem 4 can be utilized for finite-horizon-based computation or value iteration, see Appendix 8.8.
4.3 FRT Value Functions and CBFs
@@
We first derive an optimal policy of $V_{\gamma}$ from the HJ-FRT-VI (21).
@@
where $V_{\gamma}$ is defined as (18). Then, $K_{\gamma}(x)$ is non-empty for every $x\in\mathrm{FRT}(\mathrm{Int}(C))$ where $\frac{\partial V_{\gamma}}{\partial x}$ exists. In addition, if $V_{\gamma}$ is differentiable, any element of $K_{\gamma}(x)$ is an optimal control input with respect to $V_{\gamma}$ in (18).
@@
Note that the non-emptiness of $K_{\gamma}(x)$ is derived from $V_{\gamma}$ satisfying the HJ-FRT-VI (21). By noting that the second term of the minimum in (21) has to be non-negative for (21) to hold, we get that
$\max_{u\in U}\min_{d\in D}\frac{\partial V_{\gamma}}{\partial x}\cdot f(x,u,d)+\gamma V_{\gamma}(x)\geq 0\vskip-2.5pt$ (25)
@@
Since the value function is Lipschitz continuous and differentiable almost everywhere by Proposition 4, $V_{\gamma}$ satisfies (25) almost everywhere in $\mathrm{FRT}(\mathrm{Int}(C))\subset\mathbb{R}^{n}$. Note that $V_{\gamma}$ is 0 everywhere outside $\mathrm{FRT}(\mathrm{Int}(C))$. If $V_{\gamma}$ is differentiable in the FRT, (25) is satisfied everywhere in FRT, which constitutes the robust CBF in Definition 3:
@@
If $V_{\gamma}$ is continuously differentiable in $\mathrm{FRT}(\mathrm{Int}(C))$, $V_{\gamma}$ is a robust CBF, and $\mathrm{FRT}(\mathrm{Int}(C))$ becomes a robust control invariant set.
Next, any valid robust CBF $h_{S}$ itself is the FRT value function in $\mathrm{Int}(S)$:
@@
Let $h_{S}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a robust CBF for a closed set $S$, satisfying
$\max_{u\in U}\min_{d\in D}\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)+\gamma h(x)\geq 0,\vskip-5.0pt$ (26)
@@
$V_{\gamma}(x)=\max\{0,h_{S}(x)\}\vskip-5.0pt$ (27)
is the unique viscosity solution of the HJ-FRT-VI (21) with $h_{C}(x)=h_{S}(x)$.
@@
Finally, let $W$, a continuously differentiable function, satisfy
  1. 1.
    $W(x)\geq h_{C}(x),$ (28)
  2. 2.
    $\max_{u\in U}\min_{d\in D}\nabla W(x)\cdot f(x,u,d)+\gamma W(x)\geq 0,$ (29)
or, equivalently, it satisfies the variational inequality
[…]
for all states $x$ at which $W(x)\geq 0$. Such $W$ is a supersolution of the HJ-FRT-VI (21).
Such $W$ is a valid control barrier function whose zero-superlevel set, $C_{W}:=\{x\;|\;W(x)>0\}$, is a superset of $C$. Then, we can prove that everywhere $W(x)$ is positive, $W(x)$ is lower-bounded by $V(x)$, thus resulting in $C_{W}$ to be a superset of $\mathrm{FRT}(\mathrm{Int}(C))$:
Theorem 6.
$V_{\gamma}(x)\leq W(x)$ for all $x$ such that $V_{\gamma}(x)>0$ or $W(x)>0$ and $\mathrm{FRT}(\mathrm{Int}(C))\subset C_{W}$.
Proof.
By Theorem 5, for all $x$ such that $V_{\gamma}(x)>0$ or $W(x)>0$, we get
$V_{\gamma}(x)$ $=\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}\sup_{t\in(-\infty,0]}e^{\gamma t}h_{C}(\mathrm{x}^{-}(t))$